A manufacturing company uses the FIFO method in its process costing system. In the Cutting Department...

  1. A manufacturing company uses the FIFO method in its process costing system. In the Cutting Department in June, units were 25% complete with respect to conversion in the beginning work in process inventory and 30% complete with respect to conversion in the ending work in process inventory. Other data for the department for June follow:

units

Conversion costs

Beginning work in process inventory

20,000

50,000

Units started into production during the month

145,000

180,000

Units completed and transferred out

140,000

-

Ending inventory

25,000

-

Required: Calculate the cost per equivalent unit for conversion costs?
Answer #1

Solution: $1.37

Working: Units started and completed = Units completed during the period – Units in beginning inventory = 140,000 - 20,000 = 120,000

Units in ending inventory = Units started into production during the period - Units started and completed = 145,000 - 120,000 = 25,000

FIFO method equivalent units of production

To complete beginning work in process inventory:

Conversion: (20,000 * (100%-80%)

4,000

Units started and completed

120,000

Ending work in process inventory

Conversion: (25,000*30%)

7,500

Equivalent units of production

131,500

FIFO method cost per equivalent unit

Cost added during the period

180,000

Equivalent units of production

131,500

Cost per equivalent unit (180,000 / 131,500)

1.37